Circular No. 8074

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003at IN MCG +11-20-23 Further to IAUC 7955, M. Armstrong reports his discovery of an appar-

ent supernova (mag 17.2) on unfiltered CCD images taken on Feb. 15.150 and 17.062 UT. The new object is located at α = = 16^{h}38^{m}47^{s}.81, δ = +66^{o}01′23′′.7, which is ≈ 3′′ east and 13′′ north of the center of MCG +11-20-23. Nothing was visible at this position on Armstrong’s images taken on 2002 Apr. 23 and May 18 or on Palomar Sky Survey plates.

SUPERNOVA 2003ai Further to IAUC 8071, T. Matheson et al. report that a spectrum

(range 370–750 nm) of SN 2003ai (IAUC 8068), obtained by P. Berlind on Feb. 10.38 UT with the 1.5-m telescope, shows it to be a type-Ia supernova before maximum. A weak, narrow Hα emission line superposed on the spectrum of the supernova indicates a recession velocity of 10600 km/s for the host galaxy. Using this value, the supernova expansion velocity is ∼ 10400 km/s for Si II (rest 635.5 nm). The silicon line appears to be weaker in comparison with typical type-Ia supernovae at a similar phase. The spectral-feature age of the supernova is 7 ± 2 days before maximum light.

V2275 CYGNI S. Balman and A. Yilmaz, Middle East Technical University; A. Ret-

ter, University of Sydney; T. Ak, T. Saygac, and H. Esenoglu, Istanbul University; and Z. Aslan, Turkish National Observatory, Turkish National Research Council (TUBITAK), report that R-band CCD photometry, obtained with the new Russian-Turkish 1.5-m telescope at TUBITAK National Observatory, Antalya, during 2002 June 10, 12, Oct. 1, 2, 3, 5, Dec. 1, and 22 reveals the presence of large variations in the light curve of this classical nova (IAUC 7686), with a period of either 0.463 ± 0.014 day or the 1-day alias at 0.316 ± 0.007 day. The semi-amplitude of the variations is ∼ 0.5 mag. If the 0.463-day modulation is the correct period, the system has wide eclipses. Further observations are encouraged.

COMET C/2002 V1 (NEAT) Visual m_{1} estimates (B = binoculars): Jan. 18.73 UT, 6.7 (B. H.

Granslo, Fjellhamar, Norway, 7×50 B); 30.12, 5.6 (C. S. Morris, Fillmore, CA, 10x50 binoculars); Feb. 11.76, 3.0 (J. Shanklin, Ombersley, England, 10×50 B); 12.97, 2.3 (J. Bortle, Stormville, NY, 15×70 B); 14.72, 1.5 (Granslo, 12×50 B); 16.29, −0.5 (J. Gonzalez, Asturias, Spain, 11×80 B).

2003 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green